OpenStudy (anonymous):

Simplify Problem:

OpenStudy (anonymous):

\[\sqrt{2}\div \sqrt[3]{2}\]

OpenStudy (anonymous):

It is really set up like this:

OpenStudy (anonymous):

@iGreen @AnswerMyQuestions @563blackghost

OpenStudy (anonymous):

\[\sqrt{2} = 2^{1/2}\] when dividing exponents you subtract.

OpenStudy (anonymous):

Oh no it isn't

OpenStudy (anonymous):

no thats not the answer im saying squareroot is 1/2 exponent

OpenStudy (anonymous):

Okay,

OpenStudy (anonymous):

I just don't get how to solve it when it is in the "fraction" form I guess you can call it.

OpenStudy (anonymous):

\[\sqrt[3]{2} \implies 2^{1/3}\]

OpenStudy (anonymous):

\[\large \frac{ 2^{1/2} }{ 2^{1/3} } \implies 2^{1/2-1/3}\]

OpenStudy (anonymous):

Oh okay, that makes sense, So now what happens with the other half of the equation. I know you put \[\sqrt{2}=\sqrt{2}\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

How do I subtract two exponential fractions if they have two different denominators?

OpenStudy (anonymous):

its just fractions at this point

OpenStudy (anonymous):

OKay

OpenStudy (anonymous):

how would you normally do 1/2-1/3?

OpenStudy (anonymous):

You would change the denominators??? So that they were the same. SO its 2 and 1/6

OpenStudy (anonymous):

\[2\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

yes 2^(1/6)

OpenStudy (anonymous):

\[2^{1/6}\]

OpenStudy (anonymous):

Thank you